# Littlewood, Paley and Almost-Orthogonality: a theory well ahead of its time

### v1

@@ 3. What came next?
Littlewood–Paley theory was slow to enter the mainstream canon of harmonic analysis, but its scope and influence has grown exponentially over the last nearly 100 years. Indeed, Zygmund regarded it as well ahead of its time (see the remarks in Bob Fefferman’s Preface to the 3rd edition of Trigonometric Series [Zygmund3]), and Zygmund, together with Marcinkiewicz, were among the first to recognise its extraordinary flexibility and potential. Indeed Marcinkiewicz soon proved the Fourier multiplier theorem [Marcinkiewicz1939] which bears his name, and which we briefly discussed above. With remarkable prescience he already formulated his theorem in the multi-dimensional and multi-parameter setting which we discuss below.
But beyond a few sporadic results, it was not really until the 1950’s that Littlewood–Paley theory began to emerge within harmonic analysis as a general set of interlinked ideas with an inbuilt flexibility and resilience. (The first book-level treatment was given in the 1959 edition of Zygmund’s classic Trigonometric Series [Zygmund2].) And to realise this potential it was necessary to free the theory from its origins in complex variables and give it a new identity. The first step on this path was to replace holomorphic functions by harmonic functions and Poisson integrals: this allowed, in principle, a development of the theory in higher dimensions. Littlewood–Paley theory was finally freed from all vestiges of function theory with the introduction of purely real-variable methods. This came about as a result of the hugely significant and successful theory of Calderón and Zygmund from the early 1950’s which addressed Singular Integral Operators, the primordial example of which is the Hilbert transform [^6] The ubiquity of the Hilbert transform is partly explained by the fact that it transforms a function $f$ on the real line to the function whose harmonic extension to the upper half plane is precisely the conjugate function of the harmonic extension of $f$.
@@
acting on functions in $L^{p}(\mathbb{R})$. It is beyond the remit of this article to discuss Calderón–Zygmund theory in any detail, but suffice it to say that it was quickly recognised, primarily by Stein, that the Littlewood–Paley theory fell directly under its scope, and this real-variables perspective was fully developed in Stein’s Singular Integrals and Differentiability Properties of Functions [Stein] in 1970. The principal point here is that the manifestly non-linear mapping
@@
evaluated in the Hilbert space $\ell^{2}$, and, as such, can be treated directly [^7] Strictly speaking one needs to work with mollified versions of the dyadic block operators $Q_{j}$ as below here, but the passage between the two versions is not difficult, especially within a vector-valued perspective. within the vector-valued incarnation of Calderón–Zygmund theory. [^8] Alternatively, one my make bounds on the linear operator $f\mapsto\sum_{j}\pm Q_{j}f$ uniformly in all choices of $\pm$, and then use Khintchine’s inequality to pass to the quadratic expression. At a stroke, Littlewood–Paley theory simultaneously became available in any setting in which Calderón–Zygmund theory made sense, in particular in euclidean spaces $\mathbb{R}^{n}$, and could now be based entirely on real-variable methods. This gave an enormous boost to its potential influence and range of applicability.
Euclidean spaces $\mathbb{R}^{n}$ have a much richer set of symmetries than the unit circle $\mathbb{T}$, and it quickly became apparent that the quadratic expressions of Littlewood–Paley theory should be formulated with these symmetries in mind. Thus, on $\mathbb{R}^{n}$ we have the dilation symmetries $x\mapsto tx$ for $t>0$, and to these we may associate the discrete and continuous quadratic expressions or “square functions"
@@
The Fourier transfrom is an isometry of $L^{2}$ according to Plancherel’s theorem. Construction of such $\psi$ can be achieved by means of simple partition of unity arguments. We now let
@@
One then has, as a result of Calderón–Zygmund theory,
@@
for $1<p<\infty$. The discrete and continuous versions are to be regarded as essentially interchangeable twin manifestations of the same phenomenon. In the presence of the Pythagorean identities the reverse inequalities follow from the direct ones via polarisation.
At this level of generality (and indeed lower levels), relevant quadratic expressions come in many flavours, (such as $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ referred to in Section 2), but we have suppressed discussion of these variants in the interests of clarity and brevity. Imaginative constructions of such quadratic expressions extend the application of the theory to operators whose oscillatory nature lies well outside the scope of classical Calderón–Zygmund theory, see for example [BB].
In addition to the standard one-parameter dilations, there are many other families of symmetries of euclidean space. One such is the one-parameter non-isotropic dilations $x\mapsto(t^{a_{1}}x_{1},\dots,t^{a_{n}}x_{n})$ for $t>0$ and fixed exponents $a_{1},\dots,a_{n}>0$. While these can be treated by variants of the Calderón–Zygmund theory adapted to such dilations, the case of more general multi-parameter dilations $x\mapsto(t_{1}x_{1},\dots,t_{n}x_{n})$ for independent parameters $t_{1},\dots,t_{n}>0$ is somewhat different; but some results can be relatively easily established essentially via iterating the one-dimensional results. This is the framework within which the original Marcinkiewicz multiplier theorem fits. There are also dilational actions which have a hybrid nature between these two cases. All these have been systematically studied by Ricci and Stein and their collaborators, see for example [RicciStein].
On the other hand, systematic use of Littlewood–Paley (and vector-valued) theory in the non-isotropic setting (see especially [MR837527], [Christ] and [RicciStein]) feeds back into classical singular integral theory, and leads to a self-improving structural broadening of its aegis. For example, operators on $L^{p}(\mathbb{R}^{n})$ such as the Hilbert transform along the curve $\Gamma(t)=(t^{a_{1}},\dots,t^{a_{n}})$ given by
@@
and its associated maximal function both in the euclidean and homogeneous group settings, as studied by Christ, Fabes, Nagel, Ricci, Rivière, Stein, Wainger and others, (see for example [SW], [Christ], [RicciStein]), fall under this wider scope. This is despite the fact that such operators have a singularity which is far too severe to allow direct treatment by the classical Calderón–Zygmund theory.
When one passes to non-dilational symmetries such as translations or rotations, however, the corresponding Littlewood–Paley quadratic expressions fall outside the direct purview of Calderón–Zygmund theory. They may nonetheless still contain a great deal of geometrical content relating to almost-orthogonality and be deeply relevant. See especially Section 9 below.
An example of the flexibility offered by the real-variables formalism of Littlewood–Paley theory is the ease with which it characterises whole scales of function spaces including Sobolev spaces, Bessel potential spaces, Triebel–Lizorkin spaces and Besov spaces. (Here we work with standard dilations of $\mathbb{R}^{n}$.) Indeed, the (homogeneous) Triebel–Lizorkin spaces $F^{p,q}_{s}$ – which when $q=2$ correspond to the Bessel potential and Sobolev spaces $L^{p}_{s}$, and in particular, for $s=0$, to plain vanilla $L^{p}$ – are characterised by finiteness of the expression
@@
Amusingly, another instance of taking norms in the the “wrong" order occurs below in Section 9.
Littlewood–Paley theory was now ripe to be developed to its fullest extent.
@@ 4. Quasi-abstract Littlewood–Paley theories
The first level of abstraction takes place in the realm of conditional expectation and martingales (and thus martingale differences) on probability spaces or $\sigma$-finite measure spaces. This development was to a very large extent presaged by Paley’s 1932 paper [MR1576148], and culminated in the 1960’s and 1970’s with seminal works of Burkholder, Davis and Gundy [MR208647], [MR221573] and [MR400380]. (Interestingly, Burkholder [MR208647] credits Gundy as being the first to recognise that Paley’s paper was susceptible to substantial generalisation.) The central result is the analogue of Theorem 2.1 in this setting, the equivalence of $L^{p}$ norms of a martingale with those of its martingale difference square function. This subsumed in turn an earlier result of Marcinkiewicz and Zygmund [MZ] with a similar flavour concerning independent random variables with mean zero. In conjunction with the remark after Theorem 2.1, that result explains why lacunary Fourier series are sometimes said to behave in the same way as sums of independent random variables. The Marcinkiewicz–Zygmund result in turn subsumes Khintchine’s inequality. It is the results of Burkholder, Davis and Gundy which unlock all of the potential of Littlewood–Paley theory from Fourier Analysis into the stochastic and indeed more general settings.
@@
Simultaneously with Burkholder’s work, Stein was beginning to develop Littlewood–Paley theory on compact Lie groups. He realised that the correct setting for this was that of symmetric diffusion semigroups (mimicking the Poisson semigroup in euclidean spaces). In its most general formulation, his theory relies on the Hopf–Dunford–Schwartz maximal ergodic theorem, together with estimates for quadratic expressions which are (nontrivially!) obtained from the corresponding ones in the setting of martingale differences. This, together with many other developments concerning the essentially holomorphic nature of the semigroup setting and its application to holomorphic spectral multipliers, is set out in [SteinTHA]. Later, Cowling [Cowling1], [Cowling2] offered an alternative and perhaps simpler perspective – based upon transference – on this circle of ideas.
@@
These developments – in both the martingale and the semigroup settings – have had a profound influence which is worthy of far more extensive discussion than we can give here. For now we wish merely to note that the power that abstraction brings can be exploited and re-channelled to give fresh insight into even the most classical of settings, with spectacular results, particularly when it comes to the quantitative study of bounds. Indeed, the semigroup approach leads to dimension-free bounds for the Hardy–Littlewood maximal operator defined over balls, and even over arbitrary convex bodies in $\mathbb{R}^{n}$ (see [SteinBig], [Bourgain], [CarbConvex]), and martingale methods lead to the best constant in the $\ell^{p}(\mathbb{Z})$ bounds for the discrete version of the Hilbert transform (5). In this latter context see especially the survey article of Bañuelos [ban].
@@ 5. Vector-valued analysis and weighted inequalities
Vector-valued analysis arising out of Littlewood–Paley theory brings a new flexibility of perspective. In short, instead of considering scalar-valued operators $T$, we consider operators taking possibly vector-valued inputs to possibly vector-valued outputs. The whole machinery of singular integral operators adapts readily to this situation. As we have seen above, Littlewood–Paley theory neatly fits this paradigm, (in which case the domain consists of scalar-valued functions and the codomain of Hilbert-space-valued functions). The Marcinkiewicz–Zygmund result saying that any bounded linear operator on $L^{p}$ automatically extends to a bounded linear operator on $L^{p}(\ell^{2})$ with the same norm is an early example of the vector-valued philosophy. Apart from anything else, vector-valued inequalities form an essential part of the working toolkit of modern Fourier analysis. There is nothing limiting us to considering only Hilbert spaces for our vectors, and indeed when considering general Banach spaces, properties such as UMD (unconditional martingale differences) turn out to play a key role (for example, the Hilbert transform is bounded on $L^{p}$ with values in a Banach space $X$ if and only if $X$ is UMD, see [Bourg1]).
@@
for $p>2$ such that $(p/2)^{\prime}=r$. It is a quite remarkable fact that the converse to this holds: under rather general conditions, an inequality of the sort (8) necessarily implies that some inequality of the sort (7) holds. And the principle that vector-valued inequalities are equivalent to weighted inequalities is quite general. But the catch is that the argument is non-constructive: even if we know perfectly what are the $T_{j}$, we in general have no idea what will be the mapping $w\mapsto\mathcal{M}w$. Even at the level of a single $T_{j}$ this remains remarkable, and leads to the philosophy that “there is no such thing as $L^{p}$ – only $L^{2}$ with weights".
See the book [GCRdeF] by García-Cuerva and Rubio de Francia for an extensive introduction to these topics and their interplay with Littlewood–Paley theory until the mid 1980’s. The theory has continued to develop apace since then, not least in the context of non-commutative and operator-valued harmonic analysis, building on the martingale perspective. See [LuPiqPis] and [PisierXu]. Littlewood–Paley inequalities are valid in $L^{P}(X)$ spaces if and only if $X$ is UMD, see [MR752501]. For some more recent developments see [KLW].
@@ 6. Wavelets and the Calderón reproducing formula
Thanks largely due to the incisive insights of Yves Meyer, for which he was awarded the Abel Prize, there has arisen since the 1980’s a vast literature on wavelets and their relatives, to which we cannot possibly do justice. Three excellent references are [Meyer], [Daubechies] and [HernandezWeiss]. Beyond this, the applications of wavelets across science and engineering are countless.
@@
where $D_{jk}$ is the dyadic cube of sidelength $2^{j}$ anchored at $2^{j}k$. Thus, in particular, all the information about the size of an $f\in L^{p}(\mathbb{R}^{n})$ can be read off from just the size of its wavelet coefficients; this is in stark contrast to what happens in classical Fourier analysis, where the size of the Fourier coefficients is far from the full story, and it makes computation in the wavelet setting much more stable than its Fourier counterpart.
@@
The second and third terms arising here are examples of paraproducts, so-called because they mimic classical products but with a limited interaction between Fourier frequencies.
One of the natural questions one may pose about such decompositions is whether there are estimates for each of the three terms which allow us to recover the Hölder inequality $\|fg\|_{p}\lesssim\|f\|_{q}\|g\|_{r}$ where $1/p=1/q+1/r$. The first term does not present difficulties, and can be treated by Cauchy–Schwarz and then Littlewood–Paley theory directly. The paraproduct terms can be treated using Carleson measures and fall under the general theory of non-convolution singular integrals of David, Journé and Semmes going by the name of $T1$ and $Tb$ theorems. See [SteinBig], [DJ], [DJS], [Hofmann]. A general singular integral operator in this class can be written as a sum of three terms, one of which has special cancellation properties such as $T1=T^{\ast}1=0$, and the other two of which are paraproducts.
Paraproducts and their relatives also arise in non-linear PDE (Bony’s paradifferential calculus) and are used to establish versions of the Leibniz formula (product rule for derivatives) for fractional derivatives.

### v2

@@ 3. What came next?
While Littlewood–Paley theory was slow to enter the mainstream canon of harmonic analysis, its scope and influence have grown exponentially over the last nearly 100 years. Indeed, Zygmund regarded Littlewood–Paley theory as well ahead of its time (see the remarks in Bob Fefferman’s Preface to the 3rd edition of Trigonometric Series [77]), and Zygmund, together with Marcinkiewicz, were among the first to recognise its extraordinary flexibility and potential. Indeed, Marcinkiewicz soon proved the Fourier multiplier theorem [55] which bears his name, and which we briefly discussed above. With remarkable prescience he already formulated his theorem in the multi-dimensional and multi-parameter setting which we discuss below.
But beyond a few sporadic results, it was not really until the 1950’s that Littlewood–Paley theory began to emerge within harmonic analysis as a general set of interlinked ideas with an inbuilt flexibility and resilience. (The first book-level treatment was given in the 1959 edition of Zygmund’s classic Trigonometric Series [76].) And to realise its potential, it was necessary to free the theory from its origins in complex variables, and give it a new identity. The first step on this path was to replace holomorphic functions by harmonic functions and Poisson integrals: this allowed, in principle, a development of the theory in higher dimensions. Littlewood–Paley theory was finally freed from all vestiges of complex function theory with the introduction of purely real-variable methods. This came about as a result of the hugely significant and successful theory of Calderón and Zygmund from the early 1950’s which addressed Singular Integral Operators, the primordial example of which is the Hilbert transform [^7] The ubiquity of the Hilbert transform is partly explained by the fact that it takes a function $f$ on the real line to the function whose harmonic extension to the upper half plane is precisely the conjugate function of the harmonic extension of $f$.
@@
acting on functions in $L^{p}(\mathbb{R})$. It is beyond the remit of this article to discuss Calderón–Zygmund theory in any detail, but suffice it to say that it was quickly recognised, primarily by Stein, that the Littlewood–Paley theory fell directly under its scope, and this real-variables perspective was fully developed in Stein’s Singular Integrals and Differentiability Properties of Functions [69] in 1970. The principal point here is that the manifestly non-linear mapping
@@
evaluated in the Hilbert space $\ell^{2}$, and, as such, can be treated directly [^8] Strictly speaking one needs to work with mollified versions of the dyadic block operators $Q_{j}$ as below here, but the passage between the two versions is not difficult, especially within a vector-valued framework. within the vector-valued incarnation of Calderón–Zygmund theory. [^9] Alternatively, one my make bounds on the linear operator $f\mapsto\sum_{j}\pm Q_{j}f$ uniformly in all choices of $\pm$, and then use Khintchine’s inequality to pass to the quadratic expression. At a stroke, Littlewood–Paley theory simultaneously became available in any setting in which Calderón–Zygmund theory made sense, in particular in euclidean spaces $\mathbb{R}^{n}$, and could now be based entirely on real-variable methods. This gave an enormous boost to its potential influence and range of applications.
Euclidean spaces $\mathbb{R}^{n}$ have a much richer set of symmetries than the unit circle $\mathbb{T}$, and it quickly became apparent that the quadratic expressions of Littlewood–Paley theory should be formulated with these symmetries in mind. Thus, on $\mathbb{R}^{n}$ we have the dilation symmetries $x\mapsto tx$ for $t>0$, and to them we may associate the discrete and continuous quadratic expressions or “square functions"
@@
The Fourier transform is an isometry of $L^{2}$ according to Plancherel’s theorem. Construction of such $\psi$ can be achieved by means of simple partition of unity arguments. We now let
@@
One then has, as a direct consequence of Calderón–Zygmund theory, the fundamental relations
@@
for $1<p<\infty$.
Remark.
(i) The discrete and continuous versions are to be regarded as essentially interchangeable twin manifestations of the same phenomenon.
(ii) In the presence of the Pythagorean identities the reverse inequalities follow from the direct ones via polarisation.
(iii) At this level of generality (and indeed lower levels), relevant quadratic expressions come in many flavours, (such as $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ mentioned in Section 2), but we have suppressed discussion of these variants in the interests of clarity and brevity.
(iv) Imaginative constructions of such quadratic expressions extend the application of the theory to operators whose oscillatory nature lies well outside the scope of classical Calderón–Zygmund theory, see for example [4].
In addition to the standard one-parameter dilations, there are many other families of symmetries of euclidean space. One such is the one-parameter non-isotropic dilations $x\mapsto(t^{a_{1}}x_{1},\dots,t^{a_{n}}x_{n})$ for $t>0$ and fixed exponents $a_{1},\dots,a_{n}>0$. While these can be treated by variants of the Calderón–Zygmund theory adapted to such dilations, the case of more general multi-parameter dilations $x\mapsto(t_{1}x_{1},\dots,t_{n}x_{n})$ for independent parameters $t_{1},\dots,t_{n}>0$ has a somewhat different character; but some results can be relatively easily established essentially via iterating the one-dimensional results. This is the framework within which the original Marcinkiewicz multiplier theorem fits. There are also dilational actions which have a hybrid nature, located between these two cases. All these have been thoroughly studied by Ricci and Stein and their collaborators, see for example [63].
On the other hand, systematic use of Littlewood–Paley (and vector-valued) theory in the non-isotropic setting (see especially [34], [21] and [63]) feeds back into classical singular integral theory, and, rather remarkably, leads to a self-improving structural broadening of its aegis. For example, operators on $L^{p}(\mathbb{R}^{n})$ such as the Hilbert transform along the curve $\Gamma(t)=(t^{a_{1}},\dots,t^{a_{n}})$ given by
@@
and its associated maximal function have a singularity which is far too severe to allow direct treatment by the classical non-isotropic Calderón–Zygmund theory. Nevertheless, they do fall under this expanded scope, and have been studied by Christ, Fabes, Nagel, Ricci, Rivière, Stein, Wainger and others, (see for example [68], [21], [63]), both in the euclidean and more generally the homogeneous group settings.
When one passes to non-dilational symmetries such as translations or rotations, however, the corresponding Littlewood–Paley quadratic expressions fall outside the direct purview of Calderón–Zygmund theory. They may nonetheless still contain a great deal of geometrical content relating to almost-orthogonality and be highly relevant for applications. See especially Section 9 below.
An example of the flexibility afforded by the real-variables formalism of Littlewood–Paley theory is the ease with which it characterises whole scales of function spaces including Sobolev spaces, Bessel potential spaces, Triebel–Lizorkin spaces and Besov spaces. (Here we work with standard dilations of $\mathbb{R}^{n}$.) Indeed, the (homogeneous) Triebel–Lizorkin spaces $F^{p,q}_{s}$ – which when $q=2$ correspond to the Bessel potential and Sobolev spaces $L^{p}_{s}$, and in particular, for $s=0$, to plain vanilla $L^{p}$ – are characterised by finiteness of the expression
@@
Amusingly, another instance of taking norms in the the “wrong" order is discussed below in Section 9.
With its real-variables foundation firmly in place, Littlewood–Paley theory was now ripe to be developed to its fullest extent.
@@ 4. Quasi-abstract Littlewood–Paley theories
The first level of abstraction takes place in the realm of conditional expectation and martingales (and thus martingale differences) on probability spaces or $\sigma$-finite measure spaces. This development was to a very large extent presaged by Paley’s 1932 paper [61], and culminated in the 1960’s and 1970’s with seminal works of Burkholder, Davis and Gundy [10], [39] and [9]. (Interestingly, Burkholder [10] credits Gundy as being the first to recognise that Paley’s paper was susceptible to substantial generalisation.) The central result is the analogue of Theorem 2.1 in this setting, the equivalence of $L^{p}$ norms of a martingale with those of its martingale difference square function. This subsumed an earlier result of Marcinkiewicz and Zygmund [54] with a similar flavour concerning independent random variables with mean zero. In conjunction with the remark after Theorem 2.1, that result explains why lacunary Fourier series are sometimes said to behave in the same way as sums of independent random variables. The Marcinkiewicz–Zygmund result in turn subsumes Khintchine’s inequality. It is the results of Burkholder, Davis and Gundy which unlock all of the potential of Littlewood–Paley theory from Fourier Analysis into the stochastic and indeed more general settings.
@@
Simultaneously with Burkholder’s work, Stein was beginning to develop Littlewood–Paley theory on compact Lie groups. He realised that the correct setting for this was that of symmetric diffusion semigroups (mimicking the Poisson semigroup in euclidean spaces). In its most general formulation, his theory relies on the Hopf–Dunford–Schwartz maximal ergodic theorem, together with estimates for quadratic expressions which are (nontrivially!) obtained from the corresponding ones in the setting of martingale differences. This, together with many other developments concerning the essentially holomorphic nature of the semigroup setting and its application to holomorphic spectral multipliers, is set out in [70]. Later, Cowling [24], [25] offered an alternative and perhaps simpler perspective – based upon transference – on this circle of ideas.
@@
These developments – in both the martingale and the semigroup settings – have had a profound influence which is worthy of far more extensive discussion than we can give here. For now we wish merely to note that the power that abstraction brings can be exploited and re-channelled to give fresh insight into even the most classical of settings, with spectacular results, particularly when it comes to the quantitative study of bounds. Indeed, the semigroup approach leads to dimension-free bounds for the Hardy–Littlewood maximal operator defined over balls, and even over arbitrary convex bodies in $\mathbb{R}^{n}$ (see [71], [6], [15]), and martingale methods lead to the best constant in the $\ell^{p}(\mathbb{Z})$ bounds for the discrete version of the Hilbert transform (5). In this latter context see especially the survey article of Bañuelos [2].
@@ 5. Vector-valued analysis and weighted inequalities
Vector-valued analysis arising out of Littlewood–Paley theory brings a new flexibility of perspective. In short, instead of considering scalar-valued operators $T$, we consider operators taking possibly vector-valued inputs to possibly vector-valued outputs. The whole machinery of singular integral operators adapts readily to this situation. As we have seen above, Littlewood–Paley theory neatly fits this paradigm, (in which case the domain consists of scalar-valued functions and the codomain of Hilbert space-valued functions). The Marcinkiewicz–Zygmund result saying that any bounded linear operator on $L^{p}$ automatically extends to a bounded linear operator on $L^{p}(\ell^{2})$ with the same norm is an early example of the vector-valued philosophy. Apart from anything else, vector-valued inequalities now form an essential part of the working toolkit of modern Fourier analysis. There is nothing limiting us to considering only Hilbert spaces for our vectors, and indeed when considering general Banach spaces, properties such as UMD (unconditional martingale differences) turn out to play a key role (for example, the Hilbert transform is bounded on $L^{p}$ with values in a Banach space $X$ if and only if $X$ is UMD, see [5], and Littlewood–Paley inequalities are valid in $L^{P}(X)$ spaces if and only if $X$ is UMD, see [57]).
@@
for $p>2$ such that $(p/2)^{\prime}=r$. It is a quite remarkable fact that the converse to this holds: under rather general conditions, an inequality of the sort (8) necessarily implies that some inequality of the sort (7) holds. This principle – that vector-valued inequalities are equivalent to weighted inequalities – is quite general. But the catch is that the argument is non-constructive: even if we know perfectly what are the $T_{j}$, we in general have no idea what will be the mapping $w\mapsto\mathcal{M}w$. Even at the level of a single $T_{j}$ this remains remarkable, and can be captured in the maxim that “there is no such thing as $L^{p}$ – only $L^{2}$ with weights".
The book [37] by García-Cuerva and Rubio de Francia contains an extensive introduction to these topics and their interplay with Littlewood–Paley theory, covering up to the mid-1980’s. The theory has continued to develop apace since then, not least in the context of non-commutative and operator-valued harmonic analysis, building on the martingale perspective. See [53] and [62], and, for some more recent developments, [46].
@@ 6. Wavelets and the Calderón reproducing formula
Thanks largely due to the incisive insights of Yves Meyer, for which he was awarded the Abel Prize, there has arisen since the 1980’s a vast literature on wavelets and their relatives, to which we cannot possibly do justice. Three excellent references are [58], [26] and [42]. Beyond this, the applications of wavelets across science and engineering are countless.
@@
where $D_{jk}$ is the dyadic cube of sidelength $2^{j}$ anchored at $2^{j}k$. Thus, in particular, all the information about the size of an $f\in L^{p}(\mathbb{R}^{n})$ can be read off from just the size of its wavelet coefficients: this is in stark contrast to what happens in classical Fourier analysis, where the size of the Fourier coefficients is far from the full story; moreover it makes computation in the wavelet setting much more stable than in its Fourier counterpart.
@@
The second and third terms arising on the right-hand side here are examples of paraproducts, so-called because they mimic classical products but with a limited interaction between Fourier frequencies.
One of many natural that questions one may pose about such decompositions is whether there are estimates for each of the three terms which allow us to recover the Hölder inequality $\|fg\|_{p}\lesssim\|f\|_{q}\|g\|_{r}$ where $1/p=1/q+1/r$. The first term does not present difficulties, and can be treated by Cauchy–Schwarz and then Littlewood–Paley theory directly. The paraproduct terms can be treated using Carleson measures and fall under the general theory of non-convolution singular integrals of David, Journé and Semmes going by the name of $T1$ and $Tb$ theorems. See [71], [29], [27], [43]. Notably, a general singular integral operator of the David–Journé–Semmes class can be written as a sum of three terms, one of which has special cancellation properties such as $T1=T^{\ast}1=0$, and the other two of which are paraproducts.
Paraproducts and their relatives also arise in non-linear PDE (in Bony’s paradifferential calculus) and are used to establish versions of the Leibniz formula (i.e. the product rule) for fractional derivatives.